OpenStudy (anonymous):

how do you factor 4asquared-4ab+bsquared-csquared? PLEASE HELP!!!!!!!

OpenStudy (inkyvoyd):

4a^2-4ab+b^2-c^2 =(2a-b)^2-c^2

OpenStudy (anonymous):

how did you do that?

OpenStudy (inkyvoyd):

remember the difference of squares formula.

OpenStudy (anonymous):

xsquared-ysquared=(x+y)(x-y)??

OpenStudy (inkyvoyd):

yes.

OpenStudy (inkyvoyd):

(2a-b)^2-c^2 =(2a-b+c)(2a-b-c)

OpenStudy (anonymous):

how does that apply?? theres more than 2 squared terms.

OpenStudy (inkyvoyd):

well, it does. I could show it to you by multiplying out, but let's do it an easier way.

OpenStudy (anonymous):

ok

OpenStudy (inkyvoyd):

(2a-b)^2-c^2 ->replace 2a-b with u. u^2-c^2=(u+c)(u-c) Replace u with 2a-b (2a-b+c)(2a-b-c)

OpenStudy (inkyvoyd):

notice how it only works if the whole thing is in parenthesis, like (2a-b)^2, not (2a)^2-b, or (2a-b^2), neither of those work. But do notice, that something like (x^4-1) can be factored as (x^2-1)(x^2+1)=(x+1)(x-1)(x^2+1)

OpenStudy (anonymous):

O i see! thank you!

OpenStudy (inkyvoyd):

No problem :)